homework question help
A barge moves 11 km/h in still water. It travels 50 km up river and 50km down river in a total of 22 h. What is the speed of the current

Answer by jorel1380(3719)

50/(11+c)+50/(11-c)=22
50(11-c)+50(11+c)=22(121-c2)
550-50c+550+50c=2662-22c2
22c2=2662-1100
22c2=1562
c2=71
c=√71=8.43
The current is 8.43 kph..